Grade 9 math question;
Here is a recipe for apple pie:
500ml flour
200ml butter
500g sliced apples
125g sugar
Each pie serves 6 people. I only have 350 g of sliced apples because Frank ate 3 apples. How much of each other ingredients will I need to make the pie proportionally correct?
Thanks :)

Answer by rfer(16322)

350/500=0.7 or 70%
500*0.7=350 flour
200*0.7=140 butter
125*0.7=87.5 sugar